A mechanic is working on a car with a 20-qt radiator containing a 60% antifreeze solution. HOw much of the solution should he drain and replace with pure water to get a solution that is 50% antifreeze.
Answer by checkley71(8403)

.6(20-X)=.5*20
12-.6X=10
-.6X=10-12
-.6X=-2
X=-2/-.6
X=3.33 QT NEEDS TO BE DRAINED & REPLACED WITH PURE WATER.
PROOF
.6(20-3.33)=.5*20
.6*16.67=10
10=10